Victor purchased shares of cisco stock as follows: 50sh for 1,550in 2012= 31.00 100sh for 2,500 in 2014=25.00 125sh for 3,500 in 2016=28.00 In 2018, he sold 150 shares at 45.00 per sh for gross proceeds of 6,750. If he uses the specific identification method to reduce or eliminate the month of capital gain from this sale, the basis he will use to compute his capital gain or loss is

I think the answer is 4,350 ?
Answer #1

Capital gain = sale proceeds - tax basis.

In the specific identification method, the shares with the highest cost will be chosen so that the capital gain can be reduced.

150 shares are sold. Therefore, 150 shares with the highest cost should be chosen.

50 shares bought in 2012 @ 31.00 per share have the highest cost. Hence these will be chosen.

Remaining shares to choose = 150 - 50 = 100.

125 shares bought in 2016 @ 28.00 per share have the next highest cost. Hence, 100 shares will be chosen from these.

Tax basis for 150 shares = (50 * 31.00) + (100 * 28.00) = 4350.